OpenStudy (anonymous):

Assume that the box contains 11 balls: 1 red, 4 green, and 6 blue. Two balls are drawn at random, one after the other without replacement.

OpenStudy (anonymous):

Is there a specific question?

OpenStudy (anonymous):

The first draw will be from 11 balls, the second draw will be out of 10 balls (since there is no replacement). Keep that in mind for computing probabilities.

OpenStudy (anonymous):

Oh what is the probability that the first ball was blue, given that the second was red?

OpenStudy (anonymous):

blue= 6/11=55% red= 1/10= 10%